# Web Privacy based on Contextual Integrity: Measuring the Collapse of Online Contexts

### v1

@@ 6.3. JS fingerprinting between- and within-contexts
The usage of JS fingerprinting by trackers to persistently identify multiple or single context collapse was 5-10 times lower than the usage of cookie IDs, across all contexts under study. The number of participating websites connected via JS fingerprinting was 2-3 times lower than their number connected via cookie IDs. When inspecting overlap between cookie ID and JS fingerprinting usage by actors for multiple context collapse, we found 13 different trackers that use both cookie IDs and JS fingerprinting. They originated from six out of our seven studied contexts (none was found in the adult context), and their applied JS fingerprinting scripts usually travel to fewer contexts than the cookie IDs they use. On average, 259 persistent identifiers use cookie IDs exclusively to collapse multiple contexts, and 24 persistent identifiers appear in our data for the first time as they exclusively use JS for multiple context collapse. Interestingly, 62 new websites, can be now tagged as participating websites in context collapse using JS fingerprinting. These websites cover sites from all contexts, distributed almost evenly among contexts.
@@
Figure 10 visualizes multiple context collapse from the eCommerce content on Nov 3, 2024. Grey nodes and edges represent context collapse that is exclusviely happening via cookie IDs. Orange nodes and edges represent user ID diffusion by both cookie IDs and JS fingerprinting. Nodes and edges in light blue are places where there has been an exclusive use of JS fingerprinting for multiple context collapse. Ten new edges and eighteen new nodes are exclusively connected via JS fingerprinting functions in this crawl.
@@ 7. Discussion & Suggestions
Our results reveal the varying structural connections of online contexts on the Web and show the extent of single and multiple context collapse, for tens of millions of users per month, on the top-700 popular websites across seven distinct contexts. The amount and portion of trackers that engage in persistent browser identification and the participating websites that embed those trackers vary between contexts. Single context collapse, the linkage of websites from the same context, tends to behave like traditional Web tracking, with many websites within a context are connected via the dominant trackers from the advertising industry, and a few trackers cover most of the websites within each context. Multiple context collapse, however, is enabled by a fewer number of websites in each context, but by more trackers on significant portions of the enabling websites. From each context, the same percentage of third parties link user IDs further to other contexts, with most user identifiers travel to six and seven contexts.
@@
Results show how prevalent single and multiple context collapse is, and how it changes based on context and type of context collapse on the popular Web. There are many possible hypotheses for why context collapse varies, and how it can be linked to the Web personas users (Mijač et al., 2018), persona of websites’ owners (Stöver et al., 2023), and business models of websites (Sørensen and Kosta, 2019; Englehardt and Narayanan, 2016). We have not collected data to explain where various patterns of context collapse come from as it goes beyond the scope of this paper and requires a mixed-methods approach.
Our contextual measurement of Web privacy can be used to inform Mozilla’s continuous efforts to advance contextual identities for users on the Web through separating browser’s cookie and local storage to different containers (Firefox, 2024). Assigning first-party websites to different storage containers prevents trackers from accessing data associated with both, practically preventing the cookie ID-based persistent identification patterns found in this paper. Mozilla previously suggested a set of contexts for users based on their (1) personal; (2) work; (3) banking; and (4) shopping (Wiki, 2022). Instead, we call for a set of site-specific containers, per context, based on our dynamic analysis of persistent browser identification within- and between-contexts.
Per context, our results can be used to realize the required number of containers and the first-party websites to include in each container in order to prevent context collapse. We build on Hu and Sastry (2020) and further develop their ’Tangle Factor’ to realize how first-party websites are connected, within and between distinct online contexts (Hu and Sastry, 2020a). We would like to open our data analysis outputs and inform the design of browsers’ cookie storage containers by finding the vertex chromatic number and list of colored nodes of our between- and within-context network graphs. Per context, the number of colors needed for nodes such that neighboring nodes which share an edge are colored differently provides the number of required containers. The list of nodes in each color provides the list of first-party websites to include in each container. Per context, the number of required containers and the list of first-party websites to include in each container can guide browser developers on how to break users’ browser experience into storage containers and prevent the collapse of that context for the user.
We demonstrate the applicability of this approach through the two figures. In figure 11 we can see the connected websites within the LGBTQ context. The vertex chromatic number of the graph is 53, meaning that we need 53 different containers to ensure that users are not being persistently identified within the top-100 popular LGBTQ websites. The list of first-party websites to include under each container is provided in the appendix.
@@
In figure 12 we can see the connected websites from multiple contexts, based on user identifiers created in the LGBTQ context. The graph shows that in order to prevent multiple contexts collapse, we need to have eight containers and include first-party websites according to their color, as shown in the list provided in the appendix.
@@
We calculated the average number of colors (=containers) required to prevent persistent identification within each of the studied contexts. The vertex chromatic numbers for within context collapse in each context is the average number across all within context graph networks created in 28 days of data collection for each context. Our results are: Adult - 13; eCommerce - 40; Education - 52; Finance - 56; Health - 48; LGBTQ - 47; and News & Media 67. This is the number of containers, on average, required to prevent single context collapse within each context. We will need eight additional containers to prevent multiple context collapse from this context to other contexts in the study.
Importantly, splitting browser’s storage into containers will not prevent persistent identification by stateless identifiers, that take a snapshot of users’ devices and browser configurations and behavior. The JS fingerprinting methods that were studied do not store any data on users’ browsers and cannot be blocked by splitting browser’s storage to separate containers. Still, our results on JS fingerprinting can increase user awareness of the websites that are vulnerable to persistent identification through this method and render them in a more restrictive way.
@@ 8. Conclusion
The web is an essential part of our lives, constructed by an array of contexts, and serves as an enormous source for tracking individuals across their interactions in different contexts. Context collapse on the Web was found pervasive both within and across contexts, albeit to different degrees, making it impossible to maintain a fragmented identity without proper protections or meaningful regulation. This is a violation of privacy according to the theory of contextual integrity. Our study aims to empirically assess context collapse by developing a new measurement for Web privacy - the degree of within- and between-context collapse.
A key driving factor of persistent identification patterns in our analysis are third-party cookie IDs. Even though they get blocked by default across various browsers, Google, holding 65 percent of the browsers’ market share as of August 2024 (Statista, 2024), has recently declared that third-party cookies are here to stay (Spadafora, 2024). We believe our analysis can inform Chrome users, show how another popular persistent identification method - JS fingerprinting - complements cookie IDs, and help users, who are often not privacy tech experts, maintain a fragmented identity on the Web. We have provided a first glimpse on how tracking differs between and within contexts, and among different contexts, removing the curtain a little bit between individuals and websites for a better understanding of the ‘invisible contracts’ we currently have with our digital service providers and ensuring they do not undermine our natural right to privacy.
We do not declare that this is an end of our contextual investigation journey of the Web. This is just a beginning and a first modest step to empirically analyze the behavior of trackers across online contexts based on the theory of privacy as contextual integrity. Looking ahead, we plan to inspect how pervasive cookie-syncing practices and the usage of tracking pixels contribute to context collapse and how different geolocations of browsing and real-user browsing habits differently experience context collapse on the Web.
@@ A.1. Between-Context Persistent Identifiers
The difference between the mean numbers of persistent identifiers from each contexts were found to be statistically significant according to the one-sided Anova test below. The figure describes all the observations of this value across contexts over time.
@@ A.2. Between-Context Participating Websites
The difference between the mean numbers of participating websites from each contexts were found to be statistically significant according to the one-sided Anova test below. he figure describes all the observations of this value across contexts over time.
@@ A.3. Within-context persistent identifiers
The difference between the mean numbers of persistent identifiers from each contexts were found to be statistically significant according to the one-sided Anova test below. The figure describes all the observations of this value across contexts over time.
@@ A.4. Within-context participating websites
The difference between the mean numbers of participating websites from each contexts were found to be statistically significant according to the one-sided Anova test below. The figure describes all the observations of this value across contexts over time.
@@ Appendix B Appendix Tables
Table 5. SimilarWeb Websites Used for Data Collection
@@
Table 8. Top Level URLs Found as Persistent Identifiers

### v2

@@ 6.3. JS fingerprinting between- and within-contexts
The usage of JS fingerprinting by trackers to persistently identify multiple or single context collapse was 5-10 times lower than the usage of cookie IDs, across all contexts under study. The number of participating websites connected via JS fingerprinting was 2-3 times lower than their number connected via cookie IDs. When inspecting overlap between cookie ID and JS fingerprinting usage by actors for multiple context collapse, we found 13 different trackers that use both cookie IDs and JS fingerprinting. They originated from six out of our seven studied contexts (none was found in the Adult context), and their applied JS fingerprinting scripts usually travel to fewer contexts than the cookie IDs they use. On average, 259 persistent identifiers use cookie IDs exclusively to collapse multiple contexts, and 24 persistent identifiers appear in our data for the first time as they exclusively use JS for multiple context collapse. Interestingly, 62 new websites can now be tagged as participating websites in context collapse using JS fingerprinting. These websites cover sites from all contexts, distributed almost evenly among contexts.
@@
Figure 10 visualizes multiple context collapse from the eCommerce context on Nov 3, 2024. Grey nodes and edges represent context collapse that is exclusviely happening via cookie IDs. Orange nodes and edges represent user ID diffusion by both cookie IDs and JS fingerprinting. Nodes and edges in light blue are places where there has been an exclusive use of JS fingerprinting for multiple context collapse. Ten new edges and eighteen new nodes are exclusively connected via JS fingerprinting functions in this crawl.
@@ 7. Discussion & Suggestions
Our results reveal the varying structural connections of online contexts on the Web and show the extent of single and multiple context collapse, for tens of millions of users per month, on the top-700 popular websites across seven distinct contexts. The amount and portion of trackers that engage in persistent browser identification and the participating websites that embed those trackers vary between contexts. Single context collapse, the linkage of websites from the same context, tends to behave like traditional Web tracking, with many websites within a context are connected via dominant trackers from the advertising industry, and a few trackers cover most of the websites within each context. Multiple context collapse, however, is enabled by a fewer number of websites in each context, but by more trackers and on significant portions of enabling websites. From each context, the same percentage of third parties link user IDs further to other contexts, with most user identifiers travel to six and seven contexts.
@@
Results show how prevalent single and multiple context collapse are, and how it changes based on context and type of context collapse on the popular Web. There are several hypotheses for why context collapse varies between contexts, linking the collapse to the Web personas of users (Mijač et al., 2018), the persona of websites’ owners (Stöver et al., 2023), and the business models of websites (Sørensen and Kosta, 2019; Englehardt and Narayanan, 2016). We have not collected data to explain where various patterns of context collapse come from, as it goes beyond the scope of this paper and requires a mixed-methods approach.
Our contextual measurement of Web privacy can be used to inform Mozilla’s continuous efforts to advance contextual identities for users on the Web through separating browser’s cookie and local storage to different containers (Firefox, 2024). Assigning first-party websites to different storage containers prevents trackers from accessing data associated with different contexts, practically preventing cookie ID-based persistent identification patterns found in this paper. Mozilla previously suggested a set of contexts for users based on their (1) personal; (2) work; (3) banking; and (4) shopping activities (Wiki, 2022). Instead, we call for a set of site-specific containers, per context, based on our dynamic analysis of persistent browser identification within- and between-contexts.
Per context, our results can be used to realize the required number of containers and the first-party websites to include in each container in order to prevent context collapse. We build on Hu and Sastry (2020) and further develop their ’Tangle Factor’ to realize how first-party websites are connected, within and between distinct online contexts (Hu and Sastry, 2020a). We would like to open our data analysis outputs and inform the design of browsers’ cookie storage containers by finding the vertex chromatic number and list of colored nodes of our between- and within-context network graphs. Per context, the number of colors needed for nodes such that neighboring nodes which share an edge are colored differently, provides the number of required containers. The list of nodes in each color provides the list of first-party websites to include in each container. Per context, the number of required containers and the list of first-party websites to include in each container can guide browser developers on how to break users’ browsing experience into storage containers and prevent the collapse of that context for the user.
We demonstrate the applicability of our approach through the two figures. In Figure 11, we can see the connected websites within the LGBTQ context. The vertex chromatic number of the graph is 53, indicating that we need 53 different containers to ensure that users are not being persistently identified within the top-100 popular LGBTQ websites. The list of first-party websites to include under each container is provided in the appendix.
@@
In Figure 12, we can see the connected websites from multiple contexts, based on user identifiers created in the LGBTQ context. The graph shows that in order to prevent multiple contexts collapse, we need to have eight containers and include first-party websites according to their color, as shown in the appendix.
@@
We replicated the analysis for each Web context. To do so, we calculated the average number of colors (=containers) required to prevent persistent identification within each of the studied contexts. The vertex chromatic numbers for within context collapse in each context is the average number across all within context graph networks created in 28 days of data collection for each context. Our results are: Adult - 13; eCommerce - 40; Education - 52; Finance - 56; Health - 48; LGBTQ - 47; and News & Media 67. These are the numbers of containers, on average, required to prevent single context collapse within each context. We will need eight additional containers to prevent multiple context collapse from this context to other contexts in the study.
Importantly, splitting browser’s storage into containers will not prevent persistent identification by stateless identifiers. The JS fingerprinting methods that were studied do not store any data on users’ browsers and cannot be blocked by splitting browser’s storage to separate containers.
@@ 8. Conclusion
The Web is an essential part of our lives, constructed by an array of contexts, and serves as an enormous source for tracking individuals across their interactions in different contexts. Context collapse on the Web was found to be pervasive both within and across contexts, albeit to different degrees, making it impossible to maintain a fragmented identity without proper protections or meaningful regulation. This is a violation of privacy according to the theory of contextual integrity. Our study aims to empirically assess context collapse by developing a new measurement for Web privacy - the degree of within- and between-context collapse.
A key driving factor of persistent identification patterns are third-party cookie IDs. Even though they get blocked by default across various browsers, Google, holding 65 percent of the browsers’ market share as of August 2024 (Statista, 2024), has recently declared that third-party cookies are here to stay (Spadafora, 2024). We believe our analysis can inform Chrome users, show how another popular persistent identification method - JS fingerprinting - complements cookie IDs, and help users, who are often not privacy tech experts, maintain a fragmented identity on the Web. We have provided a first glimpse on how tracking differs between and within contexts, allowing for a better understanding of the ‘invisible contracts’ we currently have with our websites.
We do not declare that this is an end of our contextual investigation journey of the Web. This is just a first modest step to empirically analyze the behavior of trackers across online contexts based on the theory of privacy as contextual integrity. Looking ahead, we plan to inspect how pervasive cookie-syncing practices and the usage of tracking pixels contribute to context collapse and how different geolocations of browsing and real-user browsing habits differently experience context collapse on the Web.
@@ A.1. Between-Context Persistent Identifiers
The difference between the mean numbers of persistent identifiers from each contexts were found to be statistically significant according to the one-sided ANOVA test below. Figure 13 describes all the observations of this number, across contexts, over time, and Figure 14 shows the ANOVA test results.
@@ A.2. Between-Context Participating Websites
The difference between the mean numbers of participating websites from each contexts were found to be statistically significant according to the one-sided ANOVA test below. Figure 15 describes all the observations of this number across contexts over time, and Figure 16 shows the ANOVA test results.
@@ A.3. Within-context persistent identifiers
The difference between the mean numbers of persistent identifiers from each contexts were found to be statistically significant according to the one-sided ANOVA test below. Figure 17 describes all the observations of this value across contexts over time, and Figure 18 shows the ANOVA test results.
@@ A.4. Within-context participating websites
The difference between the mean numbers of participating websites from each contexts were found to be statistically significant according to the one-sided ANOVA test below. Figure 19 describes all the observations of this value across contexts over time, and Figure 20 shows the ANOVA test results.
@@ Appendix B Appendix Tables
Table 5. Top Popular Websites Used for Data Collection
@@
Table 8. First-party Websites that also act as Persistent Identifiers